OpenStudy (anonymous):

3m - 4n=1 3m - 2n=-1 without solving witch method you would use to solve each system; graphing,substitution, or elimination

OpenStudy (anonymous):

can anyone help?

OpenStudy (anonymous):

Do you know how to use each of those methods? Besides graphing...pretty self explanatory and cumbersome

OpenStudy (anonymous):

I know how to use the elimination method

OpenStudy (anonymous):

that's about it but I wanna make sure I pick the right one

OpenStudy (anonymous):

Well...that would help a lot with this problem if you catch my drift (๑•́ ₃ •̀๑)

OpenStudy (anonymous):

lol so its ok if I use the elimination?

OpenStudy (anonymous):

3m - 4n= 1 3m - 2n=-1 Basically...you want to subtract the 2 equations right on top of 1 another or addition in other cases (w/e) it takes to get rid of 1 of the variables

OpenStudy (anonymous):

3m-3m=0 -4n-(-2n)=-2n 1-(-1)=2 -2n=2 n=-1

OpenStudy (anonymous):

then you plug in n into 1 of the equations to solve for m

OpenStudy (anonymous):

Technically, you could also sub, but elimination would be faster in this case

OpenStudy (anonymous):

okay thank you for the help means a lot @porkbuns